# Cascaded Multi-path Shortcut Diffusion Model for Medical Image Translation

### v2

@@ 1 Introduction
On the other hand, I2I diffusion models have been recently developed and show superior performance than GANs. For general-purpose I2I with DM, [33] proposed a unified framework, Palette, which adds conditional image inputs to the previously developed Denoising Diffusion Probabilistic Model (DDPM [13]), thus enabling the I2I functionality of DDPM. To reduce the randomized initialization process and improve the stability in I2I DM, direct bridging diffusion methods have been investigated. Notably, [21] developed a Brownian Bridge Diffusion Model (BBDM) that learns the translation between two domains directly through the bidirectional diffusion process, i.e. Brownian Bridge, rather than a conditional generation process. Similarly, [23] proposed a Schrodinger Bridge I2I Diffusion Model (I2SB) that directly learns the nonlinear diffusion processes between two domains. Both had shown improved I2I performance in natural image translation tasks. Similar to I2I GANs, these DM methods have been applied in medical imaging.
@@
For example, [27] utilized DDPM to synthesize artificial histopathology images with rare cancer subtypes to mitigate data imbalance problems for medical data. [25] proposed to translate CT into MRI with conditional DDPM and score-matching models. The forward and backward diffusion processes are guided by T2 MRI. [10] proposed to perform brain PET image denoising with MRI as prior information to improve image quality. [9] utilized a contextual contained network in the DM to improve the LDCT denoising. Furthermore, 2D DMs have also been employed for 3D translation tasks, including low-count PET image denoising [45], CT reconstruction [5], and MRI super-resolution and reconstruction [20]. Direct extension to 3D DM were also explored [31]. However, there are several unique challenges of DM for I2I. First, those methods require iterating over a large number of steps in the reverse process, and most methods start the generation with pure random noise [33, 25, 10, 45, 5, 20]. This protocol not only significantly slows down the translation speed, but could also lead to diverged and sub-optimal translation results if different random noise initialization were used in the input when running multiple reverse runs (Figure 1). Even though direct bridging methods [21, 23] are translation deterministic given that no random noise input is used, they still require a large number of reverse iteration steps. Another challenge of the deterministic translation is that they also cannot generate translation uncertainty maps which is crucial for medical images, since the model’s prediction error can be used to pinpoint problem areas or give clinicians more information [36, 16, 42]. It is then a unique advantage of the stochastic sampling process of the conditional DDPM [33] to obtain the uncertainty map through running the DM repeatedly with multiple random noises [42]. Therefore, it is desirable to develop an I2I DM method that can generate high-quality converged translation results with a reduced number of required iterations, while also being able to provide translation uncertainty estimation.
@@
In the field of medical field, a wide spectrum of imaging techniques provides accurate visualization of anatomical structures of human bodies, including ultra sound, computed tomography (CT), magnetic resonance imaging (MRI), and positron emission tomography (PET). Each of them has unique advantages, and can be used for different purposes. For example, T1 MRI highlights soft tissues and is extensively applied to brain imaging. PET shows metabolic function of human and can help clinicians to locate tumors. However, post-processing is an essential step to extract accurate information from these imaging modalities for diagnosis. For example, the intrinsic imprecision of various image acquisition methods; therefore, noise reduction is necessary in improving image quality. The hybrid information from different imaging modalities often help the diagnosis; thus, translating images to another modality and generating the paired images for multiple modalities has been extensively studied [1]. For example, MR images can be translated into corresponding CT images for the purpose of PET attenuation correction. During image acquisition, sparse-measurement conditions (sparse view or limited angles) might cause challenges for accurate reconstruction, so the complementary step is usually needed before using the conventional reconstruction for image reconstruction. [22] proposed sinogram completion algorithm to address the problem in CT acquisition process. All of these tasks where our goal is to translate images from source domain to the target domain, can be considered as the instance of medical image-to-image translation problem.
With the development of deep learning algorithms in computer vision, the CNN-based generative models has been proposed to be an effective method in medical image-to-image translation. Specifically, Generative Adversarial Networks (GANs) is one of the most popular network architecture, which consists of a generator and a discriminator. The generator predicts the synthetic data from random noise, and the discriminator learn to distinguish between real training data and the synthetic data. [43] jointly train a noise reducing generator and an adversarial discriminator to denoise the low-dose cardiac CT images. With the addition of a gradient-based loss function to the pix2pix model, [29] generated CT from MRI and 7T MRI from 3T MRI. [1] cascades a series of the encoder-decoder block in the generator and combined different loss functions a new loss MedGAN framework for PET denoising and MR motion correction.
However, GAN-based models are usually task-specific, so the method might be difficult to generalize to a new image-to-image translation task without application-specific modifications. In addition, training GANs can be challenging to train due to the need to balance between the optimization of generator and discriminator (e.g find the saddle point of min-max objective). The training is therefore susceptible to non-convergence and mode collapse.
In recent years, diffusion models have outperformed other methods in generative tasks in terms of its image quality and diversity. Diffusion models iteratively add random noise to the original images until images become a gaussian noise distribution in a Markov chain, and then remove noise in the reverse process to approximate the distribution of real image. A variety of applications, including image generation, image super-resolution, image in-painting, have been accomplished with the denoising diffusion probabilistic model and score-based model [13, 39], denoted as diffusion models. [33] have developed a uniform to image-to-image translation framework Palette for above image-to-image translation tasks[33]. The cascaded diffusion model, consisting of a sequence of super-resolution diffusion models, can generate high fidelity video given a text prompt [12]. In medical imaging, DDPMs can be used to synthesize artificial histopathology images with rare cancer subtypes to mitigate data imbalance problems for medical data [27]. [25] convert CT images into corresponding MRI images with conditional DDPM and score-matching models. The forward and backward of diffusion processes are guided by T2w MRI images to guarantee the sample quality. [10] proposed to perform brain PET image denoising with MRI as prior information for improved image quality. Diffusion models have also been employed to address low-count PET image denoising for 3D whole scans with different count levels [45].
However, iterating over a large number of steps in the reverse process significantly reduces sampling speed, which is problematic for larger images. The strategy to improve sampling speed is then proposed. Denoising diffusion implicit models accelerating sampling via non-Markovian diffusion process[38]. [28] uses fewer forward passes by learning the variance of the reverse diffusion process. Another problem when applying diffusion models to medical image translation is its incapability of maintain the structural information accurately, which can be caused by the loss of structured details during the forward diffusion process and cannot be completely recovered through the learned reverse diffusion [18]. Moreover, compared with GAN, it needs no adversarial part, so the training results might be indeterministic, which is critical for medical image diagnosis. With different random starting points, the reverse process will diverge, causing inconsistencies in the translated results.
Although both GAN architecture and diffusion models have individually exhibit their capability in medical image translation tasks, the potential of combining GAN-based model and diffusion models remains largely unexplored. In this paper, we propose a novel diffusion model framework for four different medical image-to-image translation tasks, including dual energy X-ray visualization, sparse-view CT reconstruction, T1-T2 MRI conversion, and low-count PET scans denoising. The framework is consisting of two stage: the initial prediction with GAN or CNN based model, and a cascaded multi-path short cut diffusion model for further refinement. The resultant framework achieves high performance and robustness across multiple tasks.
Our contributions are summarized as follows:
  • 1.
    A new deep learning framework combining the advantages of CNN or GAN models and diffusion models is developed for a suite of medical image translation tasks. The added diffusion model refines the preliminary results of CNN or GAN models. In addition, with the prior image obtained from CNN or GAN model, it reduces the training costs for the diffusion model and increase its training stability.
  • 2.
    A novel cascaded and multi-path sampling strategy in the reverse process of the diffusion model is introduced. With different random noises added to the CNN or GAN prior, the images are transformed in multiple paths, and the averaged results are sent to the next cascade for a new round of diffusion to further improve the performance.
  • 3.
    The quantitative and qualitative results of the proposed method and other state-of-the arts method are shown for four different medical image-to-image translation tasks. The analysis of the hyperparameters for the proposed cascaded multi-path diffusion framework is shown.
2 Related Work
Generative Models. Generative models aims to generate new data instances by learning the underlying distribution of dataset and synthesizing samples that mimicking the original data distribution. Variational Autoencoders (VAE) is one of the common methods that are used for generative tasks [19]. It is a Bayesian inference-based probablistic model consisting of a encoder to learn features and representations of data in latent space and a decoder to reconstruct the representations back into its original data space. Generative Adversarial Networks (GANs) are later introduced as the mainstream generative models, based on the interplay of a generator and a discrimator. Generative Adversarial Networks (GANs) have emerged as a prominent class of generative models, characterized by the dynamic interaction between two distinct neural networks: a generator and a discriminator. The foundational concept of GANs is predicated on the adversarial relationship between these networks. The generator network is designed to produce synthetic data instances from random noise inputs, while the discriminator network functions as a binary classifier, tasked with discerning genuine data samples from the fake ones produced by the generator.
Medical Image Translation. xxx
@@ 5 Discussion
The presented work also has limitations with several potential improvements that are important subjects of our future studies. Firstly, we only validated our method on three different modalities, and evaluations on more diverse applications could be included. Even though we framed CMDM as an image-space post-processing tool here, we believe it can be further tailored to specific translation problems. For example, we could include physic-informed modules, such as data consistency [35, 40], in CMDM which may further improve its applications in medical image reconstruction [49, 37]. Secondly, the current CMDM is implemented in a 2D fashion, while 3D is desirable in many medical image translation tasks. Theoretically, we could directly change all the networks in CMDM into 3D networks to enable 3D applications, but it may be infeasible with the current computation resources. For example, we attempted to employ the 3D CMDM with an input size of $256\times 256\times 128$ on an 80G H100 GPU, however, it cannot fit into the memory even with a single batch size. Alternatively, we could also utilize multi-view diffusion or 2.5D or memory-efficient strategies to scale CMDM into 3D [5, 45, 2] which will be extensively investigated in our future works. Thirdly, the inference speed is still considered relatively long as compared to previous methods, especially the classic CNN-based methods. While we discussed the trade-off between performance and speed in the previous paragraph, it is also desirable to maintain optimal performance with increased inference speed. Utilizing accelerated diffusion models, such as DDIM and Resshift [38, 48], in CMDM could potentially help achieve this goal. To accelerate the inference speed for time-critical clinical scenarios, such as real-time translation in intervention radiology, one could also consider alternative solutions. For example, we could also consider distilling the diffusion model knowledge into the one-step inference GAN model [17], such that GAN with diffusion model performance and real-time capability can be realized. Fourthly, in the current implementation of CMDM, we did not implement ways to monitor the first step’s image generation. If unsatisfactory results were generated in the first step, the error could propagate to the next step. However, this should be reflected on the CMDM final uncertainty map where increased uncertainty value, i.e. pixel-wise standard deviation, should be observed. On the other hand, we could also further include uncertainty estimation techniques, e.g. Monte Carlo Dropout [8], in the first step of cGAN, thus monitoring prior image generation. Lastly, CMDM requires data with paired images for training, but such data may not always be available in certain applications. Unpaired translation diffusion model strategies [34, 30] could also potentially be deployed here to mitigate this challenge. For example, one could consider using CycleGAN to generate the prior image, and then using a multi-path version of UNIT-DDPM [34] for further refinement of the prior image. This is an interesting direction to be investigated in our future works. Moreover, future works also include evaluations of how CMDM impacts the downstream clinical applications. For example, we will investigate if the CMDM-translated images provide similar lesion detection capability or radiomic features when compared to the ground truth images, thus validating the clinical values of our method.

### v3

@@ 1 Introduction
On the other hand, I2I diffusion models have been recently developed and show superior performance than GANs. For general-purpose I2I with DM, [28] proposed a unified framework, Palette, which adds conditional image inputs to the previously developed Denoising Diffusion Probabilistic Model (DDPM [12]), thus enabling the I2I functionality of DDPM. To reduce the randomized initialization process and improve the stability in I2I DM, direct bridging diffusion methods have been investigated. Notably, [18] developed a Brownian Bridge Diffusion Model (BBDM) that learns the translation between two domains directly through the bidirectional diffusion process, i.e. Brownian Bridge, rather than a conditional generation process. Similarly, [19] proposed a Schrodinger Bridge I2I Diffusion Model (I2SB) that directly learns the nonlinear diffusion processes between two domains. Both had shown improved I2I performance in natural image translation tasks. Similar to I2I GANs, these DM methods have been applied in medical imaging. For example, [23] utilized DDPM to synthesize artificial histopathology images with rare cancer subtypes to mitigate data imbalance problems for medical data. [21] proposed to translate CT into MRI with conditional DDPM and score-matching models. The forward and backward diffusion processes are guided by T2 MRI. [10] proposed to perform brain PET image denoising with MRI as prior information to improve image quality. [9] utilized a contextual contained network in the DM to improve the LDCT denoising. Furthermore, 2D DMs have also been employed for 3D translation tasks, including low-count PET image denoising [38], CT reconstruction [5], and MRI super-resolution and reconstruction [17]. Direct extension to 3D DM were also explored [26]. However, there are several unique challenges of DM for I2I. First, those methods require iterating over a large number of steps in the reverse process, and most methods start the generation with pure random noise [28, 21, 10, 38, 5, 17]. This protocol not only significantly slows down the translation speed, but could also lead to diverged and sub-optimal translation results if different random noise initialization were used in the input when running multiple reverse runs (Figure 1). Even though direct bridging methods [18, 19] are translation deterministic given that no random noise input is used, they still require a large number of reverse iteration steps. Another challenge of the deterministic translation is that they also cannot generate translation uncertainty maps which is crucial for medical images, since the model’s prediction error can be used to pinpoint problem areas or give clinicians more information [31, 15, 36]. It is then a unique advantage of the stochastic sampling process of the conditional DDPM [28] to obtain the uncertainty map through running the DM repeatedly with multiple random noises [36]. Therefore, it is desirable to develop an I2I DM method that can generate high-quality converged translation results with a reduced number of required iterations, while also being able to provide translation uncertainty estimation.
@@ 4 Discussion
The presented work also has limitations with several potential improvements that are important subjects of our future studies. Firstly, we only validated our method on three different modalities, and evaluations on more diverse applications could be included. Even though we framed CMDM as an image-space post-processing tool here, we believe it can be further tailored to specific translation problems. For example, we could include physic-informed modules, such as data consistency [30, 34], in CMDM which may further improve its applications in medical image reconstruction [42, 32]. Secondly, the current CMDM is implemented in a 2D fashion, while 3D is desirable in many medical image translation tasks. Theoretically, we could directly change all the networks in CMDM into 3D networks to enable 3D applications, but it may be infeasible with the current computation resources. For example, we attempted to employ the 3D CMDM with an input size of $256\times 256\times 128$ on an 80G H100 GPU, however, it cannot fit into the memory even with a single batch size. Alternatively, we could also utilize multi-view diffusion or 2.5D or memory-efficient strategies to scale CMDM into 3D [5, 38, 1, 2] which will be extensively investigated in our future works. Thirdly, the inference speed is still considered relatively long as compared to previous methods, especially the classic CNN-based methods. While we discussed the trade-off between performance and speed in the previous paragraph, it is also desirable to maintain optimal performance with increased inference speed. Utilizing accelerated diffusion models, such as DDIM and Resshift [33, 41], in CMDM could potentially help achieve this goal. To accelerate the inference speed for time-critical clinical scenarios, such as real-time translation in intervention radiology, one could also consider alternative solutions. For example, we could also consider distilling the diffusion model knowledge into the one-step inference GAN model [16], such that GAN with diffusion model performance and real-time capability can be realized. Fourthly, in the current implementation of CMDM, we did not implement ways to monitor the first step’s image generation. If unsatisfactory results were generated in the first step, the error could propagate to the next step. However, this should be reflected on the CMDM final uncertainty map where increased uncertainty value, i.e. pixel-wise standard deviation, should be observed. On the other hand, we could also further include uncertainty estimation techniques, e.g. Monte Carlo Dropout [8], in the first step of cGAN, thus monitoring prior image generation. Lastly, CMDM requires data with paired images for training, but such data may not always be available in certain applications. Unpaired translation diffusion model strategies [29, 25] could also potentially be deployed here to mitigate this challenge. For example, one could consider using CycleGAN to generate the prior image, and then using a multi-path version of UNIT-DDPM [29] for further refinement of the prior image. This is an interesting direction to be investigated in our future works. Moreover, future works also include evaluations of how CMDM impacts the downstream clinical applications. For example, we will investigate if the CMDM-translated images provide similar lesion detection capability or radiomic features when compared to the ground truth images, thus validating the clinical values of our method.